OpenStudy (anonymous):

evaluate 6! * 8!/7! * 4!

OpenStudy (anonymous):

8! = 8.7.6.5.4.3.2.1

OpenStudy (anonymous):

. == multiply

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

do the same for the rest

hartnn (hartnn):

8!=8*7! 6!=6*5*4! does this help ?

hartnn (hartnn):

so 8!/7! = 8 and 6!/4! = 6*5 so final answer = 8*6*5=??

OpenStudy (anonymous):

yes thanks !

OpenStudy (anonymous):

6!*8! = (6*5*4!)*(8*7!) hence \[\frac{ 6!*8! }{ 7!*4! }\ = \frac{ (6*5*4!)*(8*7!) }{ 7!*4! }\] which gives 6*5*8=240

OpenStudy (anonymous):

240!

hartnn (hartnn):

240 not 240! :P